OpenStudy (anonymous):

7. Graph the circle with center at (-3, 2), which also passes through the point (0, -2). Label the center and at least four points on the circle. Write the equation of the circle. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you have the center, but the radius is missing

jimthompson5910 (jim_thompson5910):

to find the radius, we need to find the distance from (-3,2) to (0,-2)

jimthompson5910 (jim_thompson5910):

how do we do this?

OpenStudy (anonymous):

idk can you help me finding it

jimthompson5910 (jim_thompson5910):

are you familiar with the distance formula

OpenStudy (anonymous):

not really

jimthompson5910 (jim_thompson5910):

the distance formula is this d = sqrt((x2-x1)^2+(y2-y1)^2) look familiar?

OpenStudy (anonymous):

a little bit :)

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

The two generic points are (x1, y1) and (x2, y2)

OpenStudy (anonymous):

so what do we plug in?

jimthompson5910 (jim_thompson5910):

In this case, the two points are (-3, 2) and (0, -2)

jimthompson5910 (jim_thompson5910):

Which means... (x1, y1) = (-3, 2) ----> x1=-3 and y1= 2 (x2, y2) = (0, -2) ----> x2=0 and y2= -2 So in short, x1=-3 y1= 2 x2=0 y2= -2

jimthompson5910 (jim_thompson5910):

Plug these values in to get d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((0-(-3))^2+(-2-2)^2) d = sqrt((0+3)^2+( -2- 2)^2) d = sqrt((3)^2+(-4)^2) And I'll let you finish

OpenStudy (anonymous):

(x2-x1)^2+(y2-y1)^2) 0-(-3)^2+-2-2^2

jimthompson5910 (jim_thompson5910):

good (sorta)

jimthompson5910 (jim_thompson5910):

you keep going to get d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((0-(-3))^2+(-2-2)^2) d = sqrt((0+3)^2+( -2- 2)^2) d = sqrt((3)^2+(-4)^2) d = sqrt(9+16) d = ???

OpenStudy (anonymous):

sqrt 25 sqrt5 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

once you take the square root of 25, you will get 5 (and the square root will go away)

jimthompson5910 (jim_thompson5910):

in other words \[\Large \sqrt{25} = 5\]

jimthompson5910 (jim_thompson5910):

so d = 5

jimthompson5910 (jim_thompson5910):

which means that the distance from (-3, 2) to (0, -2) is 5 units

jimthompson5910 (jim_thompson5910):

making the radius to be 5 units so r = 5

OpenStudy (anonymous):

so whats the center?

jimthompson5910 (jim_thompson5910):

it's given above

jimthompson5910 (jim_thompson5910):

at the very top

OpenStudy (anonymous):

-3,2?

jimthompson5910 (jim_thompson5910):

(-3,2) yes

jimthompson5910 (jim_thompson5910):

use that with the radius to find the equation